(a) The mean and standard deviation of a set of values are 25 and 5, respectively. If a constant value 5 is added to each value, the coefficient of variation of the new set of values is equal to 10%.

(b) If (A) = 90, (AB) = 40, N = 150 and (β) = 80 then (αβ) = 30.


1
Expert's answer
2022-06-13T17:56:17-0400

(a)

The mean and standard deviation of a set of values are 25 and 5, respectively.

Adding a constant "c" to each value mean will be changed by adding a constant "c" and standard deviation remain the same.

Then "\mu_{new}=20+5=25, \sigma_{new}=5."


"coefficient\ of\ variation=\dfrac{\sigma_{new}}{\mu_{new}}\cdot100\%"

"=\dfrac{5}{25}\cdot100\%=20\%\not=10\%"

The statement that the coefficient of variation of the new set of values is equal to 10% is False.


(b)

In a 2 × 2 contingency table of two attributes :


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 A\darr\setminus B\to & B & \beta & Total \\ 
\hline
 A & (AB)=40 & (A\beta)=50 & 90 \\
 \hdashline
 \alpha & (\alpha B)=30 & (\alpha \beta)=30 & 60 \\
 \hdashline
 Total & (B)=70 & (\beta)=80 & 150 & \\
 \hdashline
\end{array}"